I need 150 gallons of 10% solution but all i have is 5% and 20% solution. How much of the 5% and 20% solution do i need in order ti create 150 gallons of 10% solution?
Answer by Alan3354(69443)

I need 150 gallons of 10% solution but all i have is 5% and 20% solution. How much of the 5% and 20% solution do i need in order ti create 150 gallons of 10% solution?
===============
t = amount of 20%
f = amount of 5%
==========
t + f = 150
20t + 5f = 10*150